# Contrastive Forward-Forward: A Training Algorithm of Vision Transformer

### v1

@@ Abstract
Although backpropagation is widely accepted as a training algorithm for artificial neural networks, researchers are always looking for inspiration from the brain to find ways with potentially better performance. Forward-Forward is a new training algorithm that is more similar to what occurs in the brain, although there is a significant performance gap compared to backpropagation. In the Forward-Forward algorithm, the loss functions are placed after each layer, and the updating of a layer is done using two local forward passes and one local backward pass. Forward-Forward is in its early stages and has been designed and evaluated on simple multi-layer perceptron networks to solve image classification tasks. In this work, we have extended the use of this algorithm to a more complex and modern network, namely the Vision Transformer. Inspired by insights from contrastive learning, we have attempted to revise this algorithm, leading to the introduction of Contrastive Forward-Forward. Experimental results show that our proposed algorithm performs significantly better than the baseline Forward-Forward leading to an increase of up to 10% in accuracy and boosting the convergence speed by 5 to 20 times on Vision Transformer. Furthermore, if we take Cross Entropy as the baseline loss function in backpropagation, it will be demonstrated that the proposed modifications to the baseline Forward-Forward reduce its performance gap compared to backpropagation on Vision Transformer, and even outperforms it in certain conditions, such as inaccurate supervision [^1] This manuscript is currently under review for publication in Neural Networks. The code associated with this work will be publicly available upon acceptance in the journal: github.com/HosseinAghagol/ContrastiveFF.
@@ 2 Related Work
The origin of the idea of contrastive learning can be attributed to [27], where a Siamese network was used based on a distance metric between two representations generated from images of signatures, although at that time was not referred to contrastive learning. Later, in [15], contrastive loss was used for dimensionality reduction. However, the major popularity of contrastive learning has come from pre-training networks in a self-supervised manner [17, 28, 29, 18]. In [37, 38], the idea was introduced that in an unsupervised task, although the labels of images are unknown, knowing that an image and its augmented version belong to the same class can allow the network to be trained in a self-supervised way. Subsequently, such developments, along with the introduction of a generalization of triplet loss [25] to utilize all non-matching samples [16], resulted in the highly popular method named SimCLR [17]. SimCLR is a simple, straightforward, and effective approach that trains the network in a self-supervised manner by applying contrastive loss in the representation space. This method proposes two forward passes through the network, where the inputs for each of these forward passes are different augmentations of the same batch. In this case, the corresponding images between these two forward passes are considered positive pairs, and the rest are considered negative pairs. In [19], this method was even extended to the post-pretraining phase in a supervised manner, called SCL. In SCL, contrastive loss was adapted to a supervised version, where positive samples defined same class samples, and negative samples defined non-same class samples.
@@
In [30], by symmetrizing the original FF loss function from the point of view of positive and negative samples, the convergence speed of training was increased. In [31], a mechanism was introduced to add collaboration between layers in FF. Although this improves the accuracy of the model, it eliminates the important property of FF, which is the independence of layers during the backward stages. Along with FF, [32] presented a method based on using Euclidean distance directly as local loss functions, grounded in contrastive learning in a self-supervised manner, which is used for pretraining the network. In [33], the approach was used contrastive learning in a supervised manner by employing cosine similarity directly as the loss functions for CNNs. The use of this method within the framework of FF has been limited to pretraining before the network is fine-tuned using the BP on specific application. From this viewpoint, this idea does not present a purely FF-based method and does not achieve satisfactory performance without the BP stage. Studies [34, 35] also attempted to extend the FF to its application in CNNs. [34] focused on the effect of various CNN elements on FF, such as filter size, by evaluating only on the MNIST [7] dataset. In [35], by introducing a channel-wise competitive learning framework, the authors eliminated the need for negative data in the FF algorithm by defining a new loss function. In addition, this study proposed new convolutional blocks to better align CNNs with FF. To our best knowledge, no studies have been conducted to extend the use of FF to ViT. In our previous work [36], we introduced a new approach to FF named Contrastive Forward-Forward, focusing on Multi Layer Perceptron (MLP). In this study, with the same view of our previous work, we extend it for use in ViT and provide a more comprehensive analysis.
@@ 3.1.1 Overview of Forward-Forward
Unlike BP, which often uses a loss function at the end of the network, FF [3] proposes to place the loss function after each layer, where the predicted label is not yet available. To this end, FF proposes a new method for image classification. Initially, two new samples are generated for each image. The first sample is created by replacing the correct label, which is one-hot encoded, in the corner of the image, resulting in $x_{i}^{pos}$. The second sample is obtained by replacing the incorrect one-hot encoded label in the corner of the image, referred to as $x_{i}^{neg}$ (although, there is no insistence on this solution alone for integrating the label with the image). These two samples are then fed to the first layer through two forward passes to obtain ${f_{i}^{pos}}$ and ${f_{i}^{neg}}$. The local loss function in FF is defined as follows:
@@ 3.2 Conceptual Similarities: Forward-Forward and Supervised Contrastive Learning
In FF, each data batch is fed into the network twice: once with the correct label and once with an incorrect label, across two forward passes. In SCL, two batches that have undergone two different and random augmentation paths are also fed into the network through two forward passes. In FF, the network weights are updated such that the input of an image batch with the correct label results in higher output values from the network layers, and lower output values for batches containing images with incorrect labels. Essentially, the network is trained to create a contrast between these two scenarios. SCL also pursues the creation of contrast with a different concept, i.e. aiming to establish this contrast by promoting similarity and dissimilarity among same and non-same class samples in the representation space, respectively. These similarities between the two methods inspired us to leverage insights from contrastive learning to revise FF.
@@ 3.3 Proposed Method: Contrastive Forward-Forward
In BP with CE loss, the label is encoded with a one-hot vector and directly fed into the loss function (Fig. 2(a)). In the original FF, the one-hot vector corresponding to the label places in a corner of the image (Fig. 2(b)). Similar to [30] if we treat the label representation as a vector the same size as the image and concatenate it with the input, i.e. $x_{i}^{\text{new}}=[x_{i}^{\text{old}};s_{i}]$, the function $f$ would be defined as: $f:\mathbb{R}^{D+D}\to\mathbb{R}^{E}$. In this case, we have a label vector of length $D$ for each class, while, FF had a one-hot vector. In the positive forward pass, the label representative for $i$’th sample is $s_{i}^{pos}$, and in the negative forward pass, the label representative for $i$’th sample is $s_{i}^{neg}$, which is one of the $C-1$ other label vectors (randomly chosen). Thus, $G_{i}^{\text{pos}}$ and $G_{i}^{\text{neg}}$ will be written according to Eq. 1 as follows:
@@
For example, $G_{i}^{\text{pos}}$ can be interpreted as the similarity between the representation produced by the specified sample and the representation of its corresponding class vector. $C$ class vectors are available. In [30], these class vectors were randomly initialized and fixed. However, instead of randomly choosing values for $s$, is it not better to use the images themselves? That is, one image could be chosen as the representative for all images of the same class, and according to 5, the goodness increases when the representations are closer to this image based on dot similarity (or cosine similarity after normalization). In fact, instead of random initialization for $s$ (Fig. 2(c)), we are using one of the images (Fig. 2(d)). Following this approach, we suggest going a step further and selecting representative label randomly from the images corresponding to the class each time instead of choosing a fixed image (Fig. 2(e)). In this case, Eq. 5 can be rewritten as follows:
@@
Let set $\alpha$ to 1, this equation can be rewritten as follows:
@@ 3.4 Contrastive Forward-Forward on Vision Transformers
Fig. 3 and Algorithm 1 (left side) represent our proposed method on ViT. The operation of projecting the patched image alongside one encoder layer is treated as the first layer from the proposed method perspective. After that, each encoder layer is considered as a separate layer. In particular, if we consider that $x_{i}$ is a flattened patched sample with dimensions (#$(\#\text{patch},h\times w\times c)$ where $h$ and $w$ is the height and width of the patch and $c$ is the number of channels in the input image, the structure of the encoder (network without a classifier head) will be as follows:
[…]
Such that each encoder_layerl($input$) is defined as follows:
@@
Compared to the proposed method, Algorithm 1 (right side) shows the baseline FF training approach for ViT with two modifications compared to the original FF [3]. These modifications are: 1-concatenating the label as a patch (randomly initialized and fixed according to [30]) to the main image patches, instead of placing it as a one-hot vector and 2-feeding the mean of representation along the tokens to the loss function, which does not change the fundamental of original FF.
@@ 3.5 Marginal Contrastive Loss
The Supervised Contrastive Loss [19], in Eq. 3, is minimized when there is maximized similarity between the representations of samples from the same class. In general, the capability of developing a representation with whole encoder is higher than an individual layer of the encoder, so it is challenging to place this loss function at each layer. Since the first layers may not yet have access to high-level features, there is a potential risk of trapping in local minima. Furthermore, if the early layers create maximum proximity between samples of the same class with the features extracted up to that point, it leaves less room for adjustments in subsequent layers. To address this, we add a margin parameter to the contrastive loss function. This aims to gradually increase the closeness of representations of same-class samples from the first to the last layers. Given that $A(i)=P(i)\cup N(i)$, where $N(i)$ represents the set of indices that include non-same class samples to the $i$-th sample, Eq. 3 can be rewritten as Eq. 14.
@@
To compute the gradient of this function, consider $\tau=1$, Eq. 15 can be shown as follows:
@@
Gradient calculations in cases where $f_{k}$ is not the anchor are provided in Eqs. A.5, A.6, and A.7. Considering the update of a specific weight $w$ in the local backward pass passes through the gradient of all present representation ($\frac{\partial L_{i}}{\partial w}=\sum_{k=1}^{2B}\frac{\partial L_{i}}{\partial f_{k}}\times\frac{\partial f_{k}}{\partial w}$). Given that Eqs. 18, A.5, A.6, and A.7 (the local backward pass follows one of these equations depending on the scenario), for an arbitrary $k$, the changes in the representations generated from $r\in R(k)$ in the range $11-m<f_{k}\cdot f_{r}<$ 1 play no role in updating the weights through the representations of each other. While the entry and exit of representations into this range play a role in altering $|R(k)|$ and $|R^{c}(k)|$. This causes the method to focus more on correcting hard positive samples i.e. $f_{k}\cdot f_{r^{\prime}}<1-m$ ($r^{\prime}\in R^{c}(k)$) while allowing the model flexibility to place $11-m<f_{k}\cdot f_{r}<$ 1 ($r\in R(k)$).
@@ 4.2 Layer-wise Training Analysis
$N$ is the total number of samples, $N_{c}$ is the number of samples in class $c$, $m$ is the overall mean of the representations, $m_{c}$ is the mean of the representations within class $c$, $S_{c}$ is the within-class scatter matrix for class $c$, representing the spread of representations within each class, $S_{W}$ is the total within-class scatter matrix, representing the combined scatter within all classes and $S_{B}$ is the between-class scatter matrix, representing the variance between the means of the different classes. In this case, the Fisher criterion in the multi-class scenario is defined as follows.
@@
[…]
@@ 4.3 Improving Performance with Contrastive Forward-Forward
Our proposed method also performed better in terms of inference time (Table 2), as expected, due need for only one forward for an unseen sample, while FF requires C forward passes. However, there is a solution to match the inference time by adding second stage to the model trained with FF, SymBa and FF Collab. In this second stage, a liner projection layer is added at the end of the network, with input from outputs all layers except the first layer [3]. Then, for each sample, a label representative (a full zeros one-hot size vector in original FF or a full zero patch in our extension of FF to ViT) is input to the network along with the image. Then, only this head is trained with CE loss. Specifically, with this layer, for an unseen sample, a single forward pass is sufficient to predict the label. The result of this approach (adding the head) is also presented in Table 2, which, although achieving a desirable inference time, has significantly reduced accuracy.
Table 1: Performance comparison of Contrastive Forward-Forward (CFF & CFF+M) against Baseline Forward-Forward (FF) and other FF-based methods on various datasets with MLP and ViT networks in term of accuracy.
@@
Table 2: Performance comparison of FF and other FF based methods with or without head against proposed method on CIFAR-10 with ViT in term of accuracy (Acc.) and inference time (Infer. T) for an unseen sample. Having a head refers to a fully connected layer with SoftMax placed at the end of the network, enabling the network to predict an unseen sample in a single forward pass (corresponding to "one-pass softmax" testing procedure in [3]).
Algorithm Head Acc.(%) Infer. T (ms×$\times 10$)
@@
In Fig. 6, the number of epochs required to reach the convergence point in various scenarios is presented. In all scenarios with FF based Methods, the proposed algorithm demonstrated a faster convergence and it shows 3 to 20 times faster than baseline FF, although it is still slower than BP with CE loss function (BP+CE).

### v2

@@ Abstract
Although backpropagation is widely accepted as a training algorithm for artificial neural networks, researchers are always looking for inspiration from the brain to find ways with potentially better performance. Forward-Forward is a novel training algorithm that is more similar to what occurs in the brain, although there is a significant performance gap compared to backpropagation. In the Forward-Forward algorithm, the loss functions are placed after each layer, and the updating of a layer is done using two local forward passes and one local backward pass. Forward-Forward is in its early stages and has been designed and evaluated on simple multi-layer perceptron networks to solve image classification tasks. In this work, we have extended the use of this algorithm to a more complex and modern network, namely the Vision Transformer. Inspired by insights from contrastive learning, we have attempted to revise this algorithm, leading to the introduction of Contrastive Forward-Forward. Experimental results show that our proposed algorithm performs significantly better than the baseline Forward-Forward leading to an increase of up to 10% in accuracy and accelerating the convergence speed by 5 to 20 times. Furthermore, if we take Cross Entropy as the baseline loss function in backpropagation, it will be demonstrated that the proposed modifications to the baseline Forward-Forward reduce its performance gap compared to backpropagation on Vision Transformer, and even outperforms it in certain conditions, such as inaccurate supervision [^1] This is the preprint of our paper accepted in Neural Networks. The published version includes additional experiments (e.g., layer-wise parallel training) and is available at doi.org/10.1016/j.neunet.2025.107867.. Code is available at: github.com/HosseinAghagol/ContrastiveFF.
@@ 2 Related Work
The origin of the idea of contrastive learning can be attributed to [27], where a Siamese network was used based on a distance metric between two representations generated from images of signatures, although at that time it was not referred to contrastive learning. Later, in [15], contrastive loss was used for dimensionality reduction. However, the major popularity of contrastive learning has come from pre-training networks in a self-supervised manner [17, 28, 29, 18]. In [37, 38], the idea was introduced that in an unsupervised task, although the labels of images are unknown, knowing that an image and its augmented version belong to the same class can allow the network to be trained in a self-supervised way. Subsequently, such developments, along with the introduction of a generalization of triplet loss [25] to utilize all non-matching samples [16], resulted in the highly popular method named SimCLR [17]. SimCLR is a simple, straightforward, and effective approach that trains the network in a self-supervised manner by applying contrastive loss in the representation space. This method proposes two forward passes through the network, where the inputs for each of these forward passes are different augmentations of the same batch. In this case, the corresponding images between these two forward passes are considered positive pairs, and the rest are considered negative pairs. In [19], this method was even extended to the post-pretraining phase in a supervised manner, called SCL. In SCL, contrastive loss was adapted to a supervised version, where positive samples defined same class samples, and negative samples defined non-same class samples.
@@
In [30], by symmetrizing the original FF loss function from the point of view of positive and negative samples, the convergence speed of training was increased. In [31], a mechanism was introduced to add collaboration between layers in FF. Although this improves the accuracy of the model, it eliminates the important property of FF, which is the independence of layers during the backward stages. Along with FF, [32] presented a method based on using Euclidean distance directly as local loss functions, grounded in contrastive learning in a self-supervised manner, which is used for pretraining the network. In [33], the approach was used contrastive learning in a supervised manner by employing cosine similarity directly as the loss functions for CNNs. The use of this method within the framework of FF has been limited to pretraining before the network is fine-tuned using the BP on a specific application. From this viewpoint, this idea does not present a purely FF-based method and does not achieve satisfactory performance without the BP stage. Studies [34, 35] also attempted to extend the FF to its application in CNNs. [34] focused on the effect of various CNN elements on FF, such as filter size, by evaluating only on the MNIST [7] dataset. In [35], by introducing a channel-wise competitive learning framework, the authors eliminated the need for negative data in the FF algorithm by defining a new loss function. In addition, this study proposed new convolutional blocks to better align CNNs with FF. To our best knowledge, no studies have been conducted to extend the use of FF to ViT. In our previous work [36], we introduced a new approach to FF named Contrastive Forward-Forward, focusing on Multi Layer Perceptron (MLP). In this study, with the same view of as our previous work, we extend it for use in ViT and provide a more comprehensive analysis.
@@ 3.1.1 Overview of Forward-Forward
Unlike BP, which often uses a loss function at the end of the network, FF [3] proposes to place the loss function after each layer, where the predicted label is not yet available. To this end, FF proposes a novel method for image classification. Initially, two new samples are generated for each image. The first sample is created by replacing the correct label, which is one-hot encoded, in the corner of the image, resulting in $x_{i}^{pos}$. The second sample is obtained by replacing the incorrect one-hot encoded label in the corner of the image, referred to as $x_{i}^{neg}$ (although, there is no insistence on this solution alone for integrating the label with the image). These two samples are then fed to the first layer through two forward passes to obtain $f_{i}^{pos}$} and $f_{i}^{neg}$}. The local loss function in FF is defined as follows:
@@ 3.2 Conceptual Similarities: Forward-Forward and Supervised Contrastive Learning
In FF, each data batch is fed into the network twice: once with the correct label and once with an incorrect label, across two forward passes. In SCL, two batches that have undergone two different and random augmentation paths are also fed into the network through two forward passes. In FF, the network weights are updated such that the input of an image batch with the correct label results in higher output values from the network layers, and lower output values for batches containing images with incorrect labels. Essentially, the network is trained to create a contrast between these two scenarios. SCL also pursues the creation of contrast with a different concept, i.e. aiming to establish this contrast by promoting similarity and dissimilarity among same and non-same class samples in the representation space, respectively. These similarities between the two methods inspired us to leverage insights from contrastive learning to modify FF.
@@ 3.3 Proposed Method: Contrastive Forward-Forward
In BP with CE loss, the label is encoded with a one-hot vector and directly fed into the loss function (Fig. 2(a)). In the original FF, the one-hot vector corresponding to the label places in a corner of the image (Fig. 2(b)). Similar to [30] if we treat the label representation as a vector the same size as the image and concatenate it with the input, i.e. $x_{i}^{\text{new}}=[x_{i}^{\text{old}};s_{i}]$, the function $f$ would be defined as: $f:\mathbb{R}^{D+D}\to\mathbb{R}^{E}$. In this case, we have a label vector of length $D$ for each class, while, FF had a one-hot vector. In the positive forward pass, the label representative for $i$’th sample is $s_{i}^{pos}$, and in the negative forward pass, the label representative for $i$’th sample is $s_{i}^{neg}$, which is one of the C1C-1 other label vectors (randomly chosen). Thus, $G_{i}^{\text{pos}}$ and $G_{i}^{\text{neg}}$ will be written according to Eq. 1 as follows:
@@
For example, $G_{i}^{\text{pos}}$ can be interpreted as the similarity between the representation produced by the specified sample and the representation of its corresponding class vector. $C$ class vectors are available. In [30], these class vectors were randomly initialized and fixed. However, instead of randomly choosing values for $s$, is it not better to use the images themselves? That is, one image could be chosen as the representative for all images of the same class, and according to 5, the goodness increases when the representations are closer to this image based on dot similarity (or cosine similarity after normalization). In fact, instead of random initialization for $s$ (Fig. 2(c)), we are using one of the images (Fig. 2(d)). Following this approach, we suggest going a step further and selecting a representative label randomly from the images corresponding to the class each time instead of choosing a fixed image (Fig. 2(e)). In this case, Eq. 5 can be rewritten as follows:
@@
Let’s set $\alpha$ to 1, this equation can be rewritten as follows:
@@ 3.4 Contrastive Forward-Forward on Vision Transformers
Fig. 3 and Algorithm 1 (left side) represent our proposed method on ViT. The operation of projecting the patched image alongside one encoder layer is treated as the first layer from the proposed method perspective. After that, each encoder layer is considered as a separate layer. In particular, if we consider that $x_{i}$ is a flattened patched sample with dimensions $(\#\text{patch},h\times w\times c)$ where $h$ and $w$ are the height and width of the patch and $c$ is the number of channels in the input image, the structure of the encoder (network without a classifier head) will be as follows:
$f_{i,1}$ $=\text{encoder\_layer}_{1}(\text{MLP}_{\text{proj}}(x_{i})+E_{\text{pos}})\quad$ $l=1$
$f_{i,l}$ $=\text{encoder\_layer}_{l}(f_{i,l-1})\quad$ $l=2,3,\dots,L$
Such that each encoder_layerl(inputinput) is defined as follows:
@@
Compared to the proposed method, Algorithm 1 (right side) shows the baseline FF training approach for ViT with two modifications compared to the original FF [3]. These modifications are: 1-concatenating the label as a patch (randomly initialized and fixed according to [30]) to the main image patches, instead of placing it as a one-hot vector and 2-feeding the mean of the representation along the tokens to the loss function, which does not change the fundamentals of original FF.
@@ 3.5 Marginal Contrastive Loss
The Supervised Contrastive Loss [19], in Eq. 3, is minimized when there is maximized similarity between the representations of samples from the same class. In general, the capability of developing a representation with the whole encoder is higher than an individual layer of the encoder, so it is challenging to place this loss function at each layer. Since the first layers may not yet have access to high-level features, there is a potential risk of trapping in local minima. Furthermore, if the early layers create maximum proximity between samples of the same class with the features extracted up to that point, it leaves less room for adjustments in subsequent layers. To address this, we add a margin parameter to the contrastive loss function. This aims to gradually increase the closeness of representations of same-class samples from the first to the last layers. Given that $A(i)=P(i)\cup N(i)$, where $N(i)$ represents the set of indices that include non-same class samples to the $i$-th sample, Eq. 3 can be rewritten as Eq. 14.
@@
To compute the gradient of this function, considering $\tau=1$, Eq. 15 can be shown as follows:
@@
Gradient calculations in cases where $f_{k}$ is not the anchor are provided in Eqs. A.5, A.6, and A.7. Considering the update of a specific weight $w$ in the local backward pass passes through the gradient of all present representation ($\frac{\partial L_{i}}{\partial w}=\sum_{k=1}^{2B}\frac{\partial L_{i}}{\partial f_{k}}\times\frac{\partial f_{k}}{\partial w}$). Given that Eqs. 18, A.5, A.6, and A.7 (the local backward pass follows one of these equations depending on the scenario), for an arbitrary $k$, the changes in the representations generated from $r\in R(k)$ in the range $1-m<f_{k}\cdot f_{r}<1$ play no role in updating the weights through the representations of each other. While the entry and exit of representations into this range play a role in altering $|R(k)|$ and $|R^{c}(k)|$. This causes the method to focus more on correcting hard positive samples i.e. $f_{k}\cdot f_{r^{\prime}}<1-m$ ($r^{\prime}\in R^{c}(k)$) while allowing the model flexibility to place $1-m<f_{k}\cdot f_{r}<1$ ($r\in R(k)$).
@@ 4.2 Layer-wise Training Analysis
NN is the total number of samples, $N_{c}$ is the number of samples in class $c$, $m$ is the overall mean of the representations, $m_{c}$ is the mean of the representations within class $c$, $S_{c}$ is the within-class scatter matrix for class $c$, representing the spread of representations within each class, $S_{W}$ is the total within-class scatter matrix, representing the combined scatter within all classes and $S_{B}$ is the between-class scatter matrix, representing the variance between the means of the different classes. In this case, the Fisher criterion in the multi-class scenario is defined as follows.
@@
The operator $\text{Tr}\{.\}$ calculates the trace of a matrix. In Fig. 5, the values of this criterion for the representations generated after each layer increased from around 13 to 40. The second row of Fig. 5 shows the 2D visualizations resulting from the t-SNE transformation [47] on the representations outputted from each layer, providing a qualitative assessment that confirms the results observed with the Fisher criterion. In the next row of this figure, there are graphs that have two vertical axes. The right vertical axis in each graph corresponds to the training and validation loss values for each epoch. Although the last layers perform better in terms of the Fisher criterion and same-class sample proximity in the qualitative charts due to receiving richer features, they are faced with higher loss due to the lower margin in the loss function. In fact, this can be considered as an advantage, as the last layers are encouraged more than expected.
@@ 4.3 Improving Performance with Contrastive Forward-Forward
Our proposed method also performed better in terms of inference time (Table 2), as expected, due to the need for only one forward for an unseen sample, while FF requires C forward passes. However, there is a solution to match the inference time by adding a second stage to the model trained with FF, SymBa and FF Collab. In this second stage, a linear projection layer is added at the end of the network, with input from the outputs of all layers except the first layer [3]. Then, for each sample, a label representative (a full zeros one-hot size vector in original FF or a full zero patch in our extension of FF to ViT) is input to the network along with the image. Then, only this head is trained with CE loss. Specifically, with this layer, for an unseen sample, a single forward pass is sufficient to predict the label. The result of this approach (adding the head) is also presented in Table 2, which, although achieving a desirable inference time, has significantly reduced accuracy.
Table 1: Performance comparison of Contrastive Forward-Forward (CFF & CFF+M) against Baseline Forward-Forward (FF) and other FF-based methods on various datasets with MLP and ViT networks in terms of accuracy.
@@
$\dagger$}There is no official code available for these methods, and their implementation has been done based on the description provided in the reference papers [30, 31], and the extension of these methods for use in ViT, similar to FF, has been carried out according to our proposed approach.
Table 2: Performance comparison of FF and other $F$ based methods with or without head against proposed method on CIFAR-10 with ViT in terms of accuracy (Acc.) and inference time (Infer. T) for an unseen sample. Having a head refers to a fully connected layer with SoftMax placed at the end of the network, enabling the network to predict an unseen sample in a single forward pass (corresponding to "one-pass softmax" testing procedure in [3]).
@@
In Fig. 6, the number of epochs required to reach the convergence point in various scenarios is presented. In all scenarios with FF-based Methods, the proposed algorithm demonstrated a faster convergence and it shows 3 to 20 times faster than baseline FF, although it is still slower than BP with CE loss function (BP+CE).